OpenStudy (anonymous):

Vertical Shift y= -1+3cos(x/4-pi)

OpenStudy (jdoe0001):

\(\bf y= -1+3cos\left(\frac{x}{4}-\pi\right)\implies y= 3cos\left(\frac{x}{4}-\pi\right)-1\\ \quad \\ y = A ( Bx + C ) + D\\ \quad \\ D = \textit{vertical shift}\quad D > 0,\quad up, D < 0,\quad down\)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (jdoe0001):

yw